# Synthesizing Strong-Coupling Kohn-Luttinger Superconductivity in 2D Van der Waals materials

### v1

@@ Abstract
Introduction- Van der Waals (VdW) materials hosting two-dimensional electrons have emerged as versatile platforms for engineering and studying many exotic correlated phases of matter, including quantum Hall states [57, 71, 72], correlated Mott insulating states  [4, 73, 54, 24], excitonic insulators [70, 46, 52] , and superconductivity [5, 36, 74, 68, 61, 69, 23, 20, 61]. This field is rapidly advancing due to the unparalleled electrical tunability, broad material palette, and moiré engineering capabilities of VdW materials [49], which continue to promise the discovery of new quantum phases.
@@
In their original proposal, Kohn and Luttinger demonstrated that electron pairing can occur in metals without involving EPC, with $T_{c}/T_{F}\sim exp[-(2\ell)^{4}]$ at large $\ell$, where $T_{F}$ is the Fermi temperature [27]. While such pairing could, in principle, lead to SC, the exponentially suppressed $T_{c}$ makes experimental observation challenging. Subsequent theoretical studies have suggested that lattice effects [2, 53], van Hove singularities [22, 44, 26], the presence of multiple Fermi surfaces [10], and quantum geometry [59, 25] may enhance KL SC. In these studies, KL pairing typically favors a $d$- wave ($\ell=2$[53, 44], or $p$- wave ($\ell=1$) pairing symmetry  [53, 26, 18]. There are also studies suggest that for weakly repulsive electrons in a single-band system, pairing with $\ell<3$ is unlikely to occur [1]. These conclusions are primarily derived from weak-coupling considerations. The behaviour of KL SC at strong-coupling remains unclear to date.
@@
Figure 1: Illustration of the trilayer Hubbard model and KL electron pairing on 2D triangular lattice. (a): AAA-stacking trilayer triangular lattice with spinfull electrons($N_{l}=3$, $N_{\sigma}=2$). Within a unit cell, different Hubbard terms $U_{l\sigma,m\sigma^{{}^{\prime}}}$ are depicted. Long-range interaction $U_{mq}(r)$ terms beyond a unit cell are not shown here. (b): One of the Kohn-Luttinger second order Feynman diagrams leading to inter-layer attraction between a pair of electrons residing on the top- and bottom- layer respectively, via the polorization of the middle-layer. (c): A cartoon showing the inter-layer pairing of two electrons (solid dots) in real-space (indicated by wavy line). Here hollow dots denote holes. For clarity, only one spin species is shown.
@@
In this works, we propose that a trilaminar 2D electron system in stacked van der Waals materials, which can be described by multi-layer Hubbard models (see Fig. 1a), can be utilized to realize an inter-layer $s$- wave strong-coupling KL superconductivity. In the strong-coupling regime, where inter-layer Hubbard repulsion $U$ becomes comparable to electronic bandwidth $W$ ($U\sim W$), the effective pairing attraction $V^{*}$ exhibits a $V^{*}\propto-U$ scaling. This behavior contrasts with the conventional weak-coupling KL pairing where $V^{*}\propto-U^{2}$, and the magnetic-correlation mediated SC in cuprates where $V^{*}\propto-1/U$, $U$ are the respective repulsive coupling constants . Such a linearly growing $V^{*}$ leads to enhanced KL SC at large $U$, which can persists even in the presence of a large Coulomb repulsion $U^{*}$ between paired electrons. We further analyse the dependence of this KL SC on electron densities and long-range Coulomb repulsions. Using density functional theory (DFT) and constrained random phase approximation (cRPA) calculations, we identify several candidate 2D layered half-metals and metals, including metal-doped phosphorene and transition-metal halides [7, 34], as potential platforms for realizing this strong-coupling KL SC.
@@
Model and Methods - We consider a three-layer repulsive Hubbard model as following,
@@
where $c^{\dagger}_{il\sigma}$ ($c_{il\sigma}$) denotes the creation (annihilation) operator for a fermion at unit cell $i$, layer $l$, and with spin flavor $\sigma$. The intra-layer nearest-neighbor hopping amplitude $t$ will be used as the energy unit throughout the paper ($t\equiv 1$ ). The electron density $n_{l}=\langle n_{l}\rangle$ in each layer can be tuned by the chemical potential $\mu$ and site-energies $\epsilon_{l}$. Unless specified, we maintain identical density for all three layers: $n_{1}=n_{2}=n_{3}=n$. The Hubbard term $U_{l\sigma,m\sigma^{\prime}}$ describes repulsions between two electrons in the same unit cell with layer and spin indices ($l,\sigma$), and $(m,\sigma^{\prime})$ respectively, $(l,\sigma)\neq(m,\sigma^{\prime})$, $U_{l\sigma,m\sigma^{\prime}}\geq 0$ in our study. $U_{mq}(r)$ denotes intra- or inter-layer long-ranged Coulomb repulsions beyond a unit cell. Here, we focus on two cases on 2D triangular lattice: trilayer spinless model ($N_{l}=3,N_{\sigma}=1$) and trilayer spinfull ($N_{l}=3,N_{\sigma}=2$) model ( See Fig.1). We consider three major term of Hubbard $U_{l\sigma,m\sigma^{\prime}}$: repulsion between electrons at neighboring layers $U\equiv U_{1\sigma,2\sigma^{\prime}}\equiv U_{2\sigma,3\sigma^{\prime}}$, and $U^{*}\equiv U_{1\sigma,3\sigma^{\prime}}$ for electrons on the top- and bottom-layer. There are also on-site repulsions $U_{o}\equiv U_{l\uparrow,l\downarrow}$ in the spinfull model. The effects of non-local $U_{lm}(r)$ will be investigated at the end of the work. To solver Eq. 1, we employ determinant quantum Monte Carlo (DQMC) and dynamical mean-field theory (DMFT) [19] approaches.
@@
Figure 2: Pairing susceptibilities $\chi$ and effective pairing attractions $V^{*}$ of the spinless trilayer Hubbard model on 2D triangular lattice from $6\times 6\times 3$ DQMC simulation. (a): Pairing sueptiblities $\chi_{l}$ for $\ell=s$- wave as a function of temperature $T$ at $U=6,U^{*}=0,n_{1}=n_{2}=n_{3}=n=0.5$. The bubble contributuion $\chi^{0}_{s}$ is shown in dashed line. (b): Effective pairing susceptiblities $\chi_{\mathrm{eff}}=\chi_{l}-\chi_{0,l}$ as a function of temperature $T$. For $p$- and $d$- wave, $\chi_{\mathrm{eff}}<0$ , denoting these pairing symmetries are disfavored by correlation effects. Other parameters are the same as in (a). (c): $\chi_{s}$ and $\chi_{s}^{0}$ as functions of $U^{\prime}$ at $U=6t,n=0.5,T=0.5t$. The critical $U^{\prime}_{c}$ where $\chi_{s}-\chi_{s}^{0}\approx 0$ is used to define the effective pairing attraction $V^{*}=-U^{\prime}_{c}$. (d): $V^{*}$ as a function of $U$ at $n=0.5,T=0.5t$. (e): $V^{*}$ as a function of electron density of the middle layer $n_{2}$ at $U=4,T=0.5$. Here $n_{1}$ and $n_{3}$ are fixed as $n_{1}=n_{3}=0.5$. (f): Inverse $s$- wave pairing susceptibility as a function of temperature $T$. Here $U=6t,n=0.6$. $\chi^{-1}\rightarrow 0$ is indicated as where $T$ approaches superconducting transition temperature $T_{c}$. Here $U^{*}=0$ for all plots.
weak- and strong-coupling limits We first consider tri-layer spinless model with $U_{12}=U_{23}=U,U_{13}=U^{*}=0$, namely, assuming the Coulomb repulsion $U^{*}$ between top- and bottom- layer ($l=3$) are fully screened off. In weak-coupling limit $U\ll t$, the second order KL Feynman diagrams give rise to an inter-layer attraction between top- and bottom- layer (within a unit cell), via the polarization of the middle-layer ($l=2$), see Fig. 1b and Fig. 1c. In the high temperature limit, this effective attraction $V_{\mathrm{eff}}$ scales like (see Appendix),
$V_{\mathrm{eff}}$ $\propto$ $-U^{2}n_{2}(1-n_{2})$ (2)
@@
where $n_{2}$ is the electron filling in the middle-layer. In the large $U/t>>1$ limit at half-filling, to avoid repulsion between electrons (holes) on neighbouring layers, if the top- and bottom- layer are occupied by a pair of electrons (holes), the middle-layer will be occupied by a hole (electron). This occupation configuration has an energy gain being of the order of $U$, hence generating an effective attraction $V_{\mathrm{eff}}$ between top- and bottom- layer electrons,
@@
see Fig. 1c and Appendix. Therefore, for both weak-coupling and strong-coupling cases, effective attraction between top- and bottom- layer electrons can be generated from the many-body effects of inter-layer repulsion $U$.
@@
DQMC result - Now we use the numerical exact DQMC to study the spinless model on 2D triangular lattices with $U_{12}=U_{23}=U,U_{13}=U^{*}=0$. In Fig. 2a, we present DQMC result on the pairing susceptibility $\chi=\chi_{\ell}^{(l,m)}$ (see Appendix) for electron pairs on the top- and bottom- layer ($l=1,m=3$) with $\ell$-wave pairing symmetry. In Fig. 2a, $\chi_{s}^{(1,3)}$ is plotted for $\ell=s$ wave as a function of temperature $T$ (solid line) at $U=6,n=0.5$, where one can see that $\chi_{s}^{(1,3)}$ grows rapidly as $T$ decreases. The sign problem prevents DQMC from approaching the BKT temperature $T_{BKT}$ for a diverging pairing susceptiblity $\chi$. A pragmatic approach to tackle the pairing tendency in DQMC is to estimate the effective pairing $\chi_{\mathrm{eff}}=\chi-\chi_{0}$, where $\chi$ and $\chi_{0}$ are respectively the full pairing susceptibility and bubble contribution. The sign of $\chi_{\mathrm{eff}}$ reflects whether the vertex corrections, i.e., the interaction effects beyond single-particle level, enhance superconductivity ($\chi_{\mathrm{eff}}>0$ ) or suppress superconductivity ($\chi_{\mathrm{eff}}<0$ ). Fig. 2a and Fig. 2b suggests that the $s$-wave pairing are favoured by correlation effects as $\chi_{eff}>0$, whereas the $p$- and $d$- wave pair as disfoured since $\chi_{\mathrm{eff}}<0$ for them. We have also considered other pairing correlations, such as the $s$-,p,dp-,d- wave pairings between the top- and middle- layers [$(l,m)=(1,2)$], where they are all disfavored (data not shown).
@@
To quantify the Kohn-Luttinger (KL) pairing mechanism, we introduce an effective attraction strength, $V^{*}$ to describe the attraction between two pairing electrons. This quantity is defined by probing the system’s response to pair-breaking repulsion. Specifically, we calculate $\chi_{\mathrm{eff}}$ for $s$- wave pairing between top- and bottom- layer electrons while introducing an explicit interlayer repulsion $U^{{}^{\prime}}$ between them. For $U^{{}^{\prime}}=0$ , we observe $\chi_{\mathrm{eff}}>0$, indicative of favored pairing, like shown in Fig.2b. Increasing $U^{{}^{\prime}}$ suppresses $\chi_{\mathrm{eff}}$, and we can identify a critical $U_{c}^{{}^{\prime}}$ where $\chi_{\mathrm{eff}}\approx 0$. We then define the effective attraction $V^{*}$ as $V^{*}=-U_{c}^{{}^{\prime}}$, as shown in Fig. 2c. This quantity measures the binding strength between two electrons opposing the pair-breaking repulsion. As shown in Fig. 2d, $V^{*}$ as a function of $U$ follows a quadratic dependence $V^{*}\propto U^{2}$ for $U<5t$, consistent with the conventional KL theory [27, 55].
@@
For larger $U>5t$, the quadratic KL relation no longer holds, $V^{*}$ gradually transits to a new linear scaling, $V^{*}\propto-U$ , underscoring the fundamental difference between KL pairing at strong-couling and the conventional KL SC (dashed line in Fig. 2d). Fig. 2e shows the density dependence of $V^{*}$ at $U=6,T=0.5$, where one sees that the relation $V^{*}\propto n_{e}^{(2)}(1-n_{e}^{(2)})$ is also nicely followed. From above result, we learn that in both weak-coupling and strong-coupling limits, DQMC result of $V^{*}$ agrees well with our analytic analysis on $V_{\mathrm{eff}}$ in Eq. 6 and Eq. 10. In essence, the overall growing tendency of $V^{*}$ with $U$ suggests that one can keep increasing inter-layer Hubbard repulsion $U$ to realize a larger effective attraction $V^{*}$ between electrons, hence potentially a higher transition temperature $T_{c}$ for SC.
@@
SC instability- We now use CDMFT to compute the superconducting $T_{c}$ at small temperatures. $T_{c}$ is identified as the temperature where pairing susceptibility $\chi$ diverges in DMFT. In Fig. 2f, we plot the inverse of $s$- wave pairing susceptibility, $\chi^{-1}$ as a function of $T$ at $U=6,n=0.6$. DQMC data available at high temperatures are also presented, where excellent agreement between DMFT and DQMC results can be seen. Extrapolating DMFT result of $\chi^{-1}$ to zero, $\chi^{-1}\rightarrow 0$, we obtain a superconducting $T_{c}\approx 0.12t$ for the given parameters in this plot. Repeating the calculations of $T_{c}$ for different $U$, the $U$-dependence of $T_{c}$ is obtained in Fig. 3a. In this plot, one sees that superconducting $T_{c}$ (triangles) grows with $U$ in a broad range of $U$, in consistent with the DQMC result ( $V^{*}$ versus $U$ in Fig. 2d). Specifically, escalated $T_{c}\gtrsim 0.12t$ can be found when $U$ is in the range of $U\sim(6t,10t)$, namely, when $U$ becomes comparable with bandwidth $W$ ($W=9t$ for 2D triangular lattice), KL SC can be greatly enhanced. For comparison, in the magnetic-fluctuation mediated SC in the single-band 2D Hubbard model, the optimal superconducting $T_{c}$ obtained by CDMFT is commonly less than $T_{c}\lesssim 0.05t$[16, 38, 35]. This means that in terms of the reduced pairing energy scale, $T_{c}/t$, the strong-coupling KL SC in the trilayer Hubbard model has an intrinsically stronger pairing force than that in the magnetic-fluctuation mediated SC in Hubbard model.
@@
Figure 4: Phase diagrams showing superconducting instability (SC) and density wave (DW) instability. (a) Superconducting transition temperature $T_{c}$ as a function of electron density $n$ on the 2D triangular lattice at $U=6$. (b) Same as (a) but at $U=12$. (c) Density wave and superconducting transition temperatures as a function of electron density $n$ on the 3D cubic lattice at $U=10$. Here for all three cases $n_{1}=n_{2}=n_{3}=n$. The repulsion between layer- ( spin flavor-) 1 and layer- (spin flavor-) 3, $U_{13}=U^{{}^{\prime}}=0.$
@@
Furthermore, the nonzero $U^{*}$ cases shown in Fig. 3a (dots and diamonds) suggest that in the presence of significant remnant Coulomb repulsion ($U^{*}=U/3,U/2$ ) between paired electrons, large superconducting $T_{c}$ can be still maintained. In Fig. 3b and Fig. 3c, $T_{c}$ as a function of $U^{*}$ are shown at $n=0.6,0.7$ respectively (for fixed $U=8t$), where finite $T_{c}$ can be found even when $U^{*}\geq U/2$. Specifically, for the $n=0.6$ case, an enhanced $T_{c}\approx 0.05t$ is found at $U=8t,U^{*}=4t$. This means that in the large screening length limit $\lambda_{TF}\rightarrow\infty$ along c-axis, i.e., when Coulomb potential in the vertical direction becomes, $U(r_{z})\propto\frac{e^{-r_{z}/\lambda_{TF}}}{r_{z}}\sim\frac{1}{r_{z}}$, (thus $U^{*}\sim U/2$ if layer distances $d_{12}=d_{23}$), the strong-coupling KL SC can still survive. This finding obviously lie beyond the expectation of high-temperature DQMC results presented in Fig. 2d, where $V^{*}\sim-U/2$ for large $U$. In a purely static picture, a condition of $U^{*}\gtrsim U/2$ leads to a net repulsion between top- and bottom-layer electrons [ i.e., $(U^{*}+V^{*})\gtrsim 0$], thereby severely inhibiting pairing between them. Hence, the persistence of a high $T_{c}$ under $U^{*}\gtrsim U/2$ suggests that the retardation part of pairing force may play a role at low temperatures (see discussions in Appendix). A further key point to emphasize is the sizeable $T_{c}$ up to a filling of $n=0.7$ (40% electron doping), as evidenced in Fig. 3c. This presents a clear difference with the case in Hubbard model for cuprates [16], where SC, induced by magnetic correlations, rapidly vanishes beyond $p\lesssim 30\%$ for hole-doping and $p\lesssim 20\%$ for electron-doping. This distinctness is rooted in the “pairing glue”: whereas SC in cuprates relies on the collective spin fluctuations [13] that are suppressed as $n\gg 0.5$ or $n\ll 0.5$ , the strong-coupling KL SC identified here does not involve any intermediate collective modes, thus ensuring robustness across a broad range of $n$.
@@
The phase diagrams - Fig. 4a presents a superconducting phase diagram showing $T_{c}$ as a function of electron density $n$ at $U=6,U^{*}=0$ for the spinless model on 2D triangular lattice from nearly empty ($n\sim 0.2$) to nearly full-filled ($n\sim 0.9$) at $U=6t,U^{*}=0$. One sees that $T_{c}$ is maximumized as $n$ being close to half-filling, in agreement with the $V^{*}$ versus $n$ tendency in DQMC (Fig. 2e). Again, here we find KL SC instability in an extremely large range of $n$. Increasing Hubbard repulsion to $U=12t$ can even further expand the SC regime. For instance, on the electron doped side, KL SC extends to $n_{c}\gtrsim 0.95$ as $T\rightarrow 0$ ( Fig. 4b).

### v2

@@ Abstract
Introduction- Van der Waals (VdW) materials hosting two-dimensional electrons have emerged as versatile platforms for engineering and studying many exotic correlated phases of matter, including quantum Hall states [57, 73, 74], correlated Mott insulating states  [4, 75, 54, 23], excitonic insulators [71, 46, 52] , and superconductivity [5, 36, 76, 69, 61, 70, 22, 19, 61]. This field is rapidly advancing due to the unparalleled electrical tunability, broad material palette, and moiré engineering capabilities of VdW materials [49, 78], which continue to promise the discovery of new quantum phases.
@@
Figure 1: Illustration of the trilayer Hubbard model and KL electron pairing on 2D triangular lattice. (a): AAA-stacking trilayer triangular lattice with spinfull electrons. Different Hubbard terms $U_{l\sigma,m\sigma^{{}^{\prime}}}$ within a unit cell are depicted. Long-range interaction $U_{mq}(r)$ terms beyond a unit cell are not shown here. (b): A KL second order Feynman diagram that leads to inter-layer attraction between a pair of electrons in the top- and bottom- layer respectively, via the polorization of the middle-layer. (c): A cartoon showing the inter-layer pairing of two electrons /holes (solid/hollow dots) in real-space (indicated by wavy line). The spin label is omitted in (b) and (c) for clarity. (d): Effective pairing attraction $V^{*}$ as a function of $U$ for spinless trilayer Hubbard model from $6\times 6\times 3$ DQMC simulation at $T=0.5,n=0.6$, cf. Fig. 2.
In their original proposal, Kohn and Luttinger demonstrated that electron pairing can occur in metals without involving EPC, below $T_{c}/T_{F}\sim exp[-(2\ell)^{4}]$ at large $\ell$, where $T_{F}$ is the Fermi temperature [27]. While such pairing could, in principle, lead to SC, the exponentially suppressed $T_{c}$ makes experimental observation challenging. Subsequent theoretical studies have suggested that lattice effects [2, 53], van Hove singularities [21, 44, 25], the presence of multiple Fermi surfaces [9], and quantum geometry [59, 24] may enhance KL SC. In these studies, KL pairing typically favors a $d$- wave ($\ell=2$[53, 44], or $p$- wave ($\ell=1$) pairing symmetry  [53, 25, 17]. There are also studies suggest that for weakly repulsive electrons in a single-band system, pairing with $\ell<3$ is unlikely to occur [1]. These conclusions are primarily derived from weak-coupling considerations. The behaviour of KL SC at strong-coupling remains unclear to date.
@@
In this work, we propose that trilaminar 2D electron systems in stacked VdW materials (see Fig. 1a), can be utilized to realize an inter-layer $s$- wave KL superconductivity. In the strong-coupling regime where inter-layer repulsion $U$ becomes comparable to electronic bandwidth $W$ ($U\sim W$), we reveal a $V^{*}\propto-U$ scaling for the effective pairing attraction $V^{*}$. This behavior contrasts with the conventional weak-coupling KL pairing where $V^{*}\propto-U^{2}/t$, and the magnetic-correlation mediated SC in cuprates where $V^{*}\propto-t^{2}/U$, $U$ and $t$ are respectively the electron repulsion and hopping amplitudes. We show that such a linearly growing $V^{*}$ leads to enhanced KL SC at large $U$. We analyse the dependence of pairing on electron densities and long-range Coulomb repulsions. Using density functional theory (DFT) and constrained random phase approximation (cRPA) calculations, we identify several candidate 2D layered half-metals and metals, including transition-metal halides [26], as potential platforms for realizing this strong-coupling KL SC.
@@
Model and Methods - We consider a repulsive Hubbard model on 2D trilayer triangular lattice as following,
@@
where $c^{\dagger}_{il\sigma}$ ($c_{il\sigma}$) denotes the creation (annihilation) operator for electrons at unit cell $i$, layer $l$, and with spin flavor $\sigma$. The intra-layer nearest-neighbor hopping amplitude $t$ will be used as the energy unit throughout the paper ($t\equiv 1$ ). The electron density $n_{l}=\langle n_{l}\rangle$ in each layer can be tuned by the chemical potential $\mu$ and site-energies $\epsilon_{l}$. Unless specified, we maintain identical density for all three layers: $n_{1}=n_{2}=n_{3}=n$. The Hubbard term $U_{l\sigma,m\sigma^{\prime}}$ describes repulsions between two electrons in the same unit cell with layer and spin indices ($l,\sigma$) and $(m,\sigma^{\prime})$ respectively, $(l,\sigma)\neq(m,\sigma^{\prime})$, $U_{l\sigma,m\sigma^{\prime}}\geq 0$ in our study. $U_{lm}(r)$ denotes intra- or inter-layer long-ranged Coulomb repulsions beyond a unit cell. Here, we will study both spinless ($N_{\sigma}=1$) and spinful ($N_{\sigma}=2$) cases ( See Fig.1). We consider three major term of Hubbard $U_{l\sigma,m\sigma^{\prime}}$: repulsion between electrons at neighboring layers $U\equiv U_{1\sigma,2\sigma^{\prime}}\equiv U_{2\sigma,3\sigma^{\prime}}$, and $U^{*}\equiv U_{1\sigma,3\sigma^{\prime}}$ for electrons on the top- and bottom-layer. There are also on-site repulsions $U_{o}\equiv U_{l\uparrow,l\downarrow}$ in the spinfull case. The effects of non-local $U_{lm}(r)$ will first be neglected, and be investigated at the end of the work. To solver Eq. 1, we employ determinant quantum Monte Carlo (DQMC) and dynamical mean-field theory (DMFT) [18] approaches.
@@
Figure 2: Pairing susceptibilities $\chi$ and effective pairing attractions $V^{*}$ of the spinless trilayer Hubbard model.(a): Pairing susceptibilities $\chi_{l}$ for $\ell=s$- wave as a function of temperature $T$ at $U=6,U^{*}=0,n=0.5$. The bubble contributuion $\chi^{0}_{s}$ is shown in dashed line. (b): Effective pairing susceptibilities $\chi_{\mathrm{eff}}=\chi_{l}-\chi_{l}^{0}$ as a function of $T$. For $p$- and $d$- wave, $\chi_{\mathrm{eff}}<0$ , denoting both pairing symmetries are unfavoured. Other parameters are the same as in (a). (c): $\chi_{s}$ and $\chi_{s}^{0}$ as functions of $U^{\prime}$ at $U=6t,n=0.5,T=0.5t$. The critical $U^{\prime}_{c}$ where $\chi_{eff}\equiv\chi_{s}-\chi_{s}^{0}\approx 0$ is used to define the effective pairing attraction $V^{*}=-U^{\prime}_{c}$. (d): Inverse $s$- wave pairing susceptibility as a function of temperature $T$. Here $U=6t,n=0.6$. $\chi^{-1}\rightarrow 0$ is indicated as where $T$ approaches superconducting transition temperature $T_{c}$. Here $U^{*}=0$ and $n_{1}=n_{2}=n_{3}=n$ for all plots. Results are obtained by $6\times 6\times 3$ DQMC simulation.
weak- and strong-coupling limits - To gain insight from an analytic perspective, we begin by considering the spinless model. Without loss of generality, we set $U_{12}=U_{23}=U$, and $U_{13}=U^{*}=0,U(r)=0$, i.e., neglecting long-ranged repulsions as well as the direct repulsion between the top and bottom layers. In weak-coupling limit $U\ll t$, second order KL Feynman diagrams yield an inter-layer attraction $V_{\mathrm{eff}}$ between the top- and bottom- layers within the same unit cell, mediated by the polarization of the middle-layer ($l=2$), as illustrated in Fig. 1b. In the instantaneous limit, $V_{\mathrm{eff}}$ scales like (see End Matter),
$V_{\mathrm{eff}}$ $\propto$ $-U^{2}n_{2}(1-n_{2})/t$ (2)
@@
where $n_{2}$ is the electron density of the middle-layer. In the large $U/t>>1$ limit at half-filling, to avoid repulsion between electrons (holes) on neighbouring layers, if the top- and bottom- layer are occupied by a pair of electrons (holes), the middle-layer will be occupied by a hole (electron). This occupation configuration has an energy gain being of the order of $U$, hence generating an effective attraction $V_{\mathrm{eff}}$ between top- and bottom- layer electrons,
@@
see Fig. 1c and End Matter. Therefore, in both weak-coupling and strong-coupling regimes, an effective attraction between top- and bottom- layer electrons can be generated from the many-body effects of inter-layer repulsion $U$, which eventually leads to superconductivity in the system.
@@
DQMC result - We now use numerical exact DQMC to study the pairing attraction. We first focus on spinless model with $U_{12}=U_{23}=U,U_{13}=U^{*}=0,U(r)=0$. In Fig. 2a, we present the pairing susceptibility $\chi=\chi_{\ell}^{(l,m)}$ for inter-layer ($l=1,m=3$) $s$-wave pairing (see End Matter) at $U=6,n=0.5$. One can see that as temperature $T$ decreases, $\chi_{s}^{(1,3)}$ (solid line) grows rapidly. The notorious sign problem prevents DQMC from accessing the Berezinskii-Kosterlitz-Thouless (BKT) temperature $T_{BKT}$ for a diverging $\chi$. A pragmatic approach to tackle the pairing tendency in DQMC is to estimate the effective pairing $\chi_{\mathrm{eff}}=\chi-\chi_{0}$, where $\chi$ and $\chi_{0}$ are respectively the full pairing susceptibility and bubble contributions. The sign of $\chi_{\mathrm{eff}}$ reflects whether the vertex corrections, i.e., the interaction effects beyond single-particle level, enhance superconductivity ($\chi_{\mathrm{eff}}>0$ ) or suppress superconductivity ($\chi_{\mathrm{eff}}<0$ ). Fig. 2a and Fig. 2b suggests that the $s$-wave pairing are favoured by correlation effects , whereas the $p$- and $d$- wave are unfavoured. We have also considered other pairing correlations, such as the $s$-,p,dp-,d- wave pairings between the top- and middle- layers [$(l,m)=(1,2)$], where they are all unfavoured (data not shown).
@@
To quantify the attraction between pairing electrons, we compute a quantity $V^{*}$ in DQMC, which is defined by probing the system’s $\chi_{\mathrm{eff}}$ response to pair-breaking repulsion. Specifically, we calculate $\chi_{\mathrm{eff}}$ for $s$- wave pairing between top- and bottom- layer electrons while introducing an explicit interlayer repulsion $U^{{}^{\prime}}$ between them. For $U^{{}^{\prime}}=0$ , one finds $\chi_{\mathrm{eff}}>0$, indicative of favoured pairing, like shown in Fig.2b. Increasing $U^{{}^{\prime}}$ will suppress $\chi_{\mathrm{eff}}$, and a critical $U_{c}^{{}^{\prime}}$ can be identified when $\chi_{\mathrm{eff}}\approx 0$. We then define the effective pairing attraction $V^{*}$ as $V^{*}=-U_{c}^{{}^{\prime}}$, as shown in Fig. 2c. This quantity naturally measures the binding strength between two pairing electrons. As shown in Fig. 1d for the spinless model, our DQMC result of $V^{*}$ exhibits a quadratic $U$ dependence, $V^{*}\propto U^{2}/t$ for small $U$ ($U\lesssim 5t$), in nice agreement with Eq. 2 that complies with the conventional KL theory [27, 55] (see also End Matter).
@@
For larger $U\gtrsim 5t$, the quadratic KL relation no longer strictly holds, $V^{*}$ gradually transits to a new linear scaling, $V^{*}\propto-U$ (dashed line in Fig. 1d), underscoring the fundamental difference between KL pairings at strong-couling and that at weak-coupling. Above results indicate that in both large and small $U$ limits, DQMC result of $V^{*}$ agrees well with our analytic analysis on $V_{\mathrm{eff}}$ in Eq. 2 and Eq. 3. In essence, the overall growing tendency of $V^{*}$ with $U$ suggests that one can keep increasing inter-layer Hubbard repulsion $U$ to realize a larger $V^{*}$ between electrons, hence potentially a higher $T_{c}$ of SC.
@@
SC instability- We now use CDMFT to compute the superconducting $T_{c}$ at low temperatures. $T_{c}$ is identified as the temperature where pairing susceptibility $\chi$ diverges in CDMFT. In Fig. 2d, we plot the inverse of $s$- wave pairing susceptibility, $\chi^{-1}$ as a function of $T$ at $U=6,n=0.6$. DQMC data available at high temperatures are also presented, where excellent agreement between DMFT and DQMC results can be seen. Extrapolating CDMFT result of $\chi^{-1}$ to zero, $\chi^{-1}\rightarrow 0$, we obtain a superconducting $T_{c}\approx 0.12t$ for the given parameters in this plot. Repeating the calculations of $T_{c}$ for different $U$, the $U$-dependence of $T_{c}$ is obtained in Fig. 3a. In this plot, one sees that $T_{c}$ (triangles) grows with $U$ in a broad parameter range, in consistent with the growing tendency of $V^{*}$ in DQMC result ( Fig. 1d). Specifically, escalated $T_{c}\gtrsim 0.12t$ can be found when $U$ is in the range of $U\sim(6t,10t)$, namely, when $U$ becomes comparable with bandwidth $W$ ($W=9t$ for 2D triangular lattice), KL SC can be greatly enhanced. For comparison, in the magnetic-fluctuation mediated SC in the single-band 2D Hubbard model, the optimal superconducting $T_{c}$ obtained by CDMFT is commonly less than $T_{c}\lesssim 0.05t$[15, 38, 35]. This means that in terms of the reduced pairing energy scale, $T_{c}/t$, the strong-coupling KL SC in the trilayer Hubbard model has an intrinsically stronger pairing force than that in the magnetic-fluctuation mediated SC in Hubbard model.
@@
Figure 4: Phase diagrams showing superconducting instability (SC) and density wave (DW) instability. (a) Superconducting transition temperature $T_{c}$ as a function of electron density $n$ on the 2D triangular lattice at $U=6t$. (b) Same as (a) but at $U=12t$. (c) Density wave and superconducting transition temperatures as a function of electron density $n$ on the 3D cubic lattice at $U=10t$. Here for all three cases $n_{1}=n_{2}=n_{3}=n$. The repulsion between layer- (pseudo-spin ) 1 and layer- (pseudo-spin) 3, $U^{*}=0.$
@@
Furthermore, the nonzero $U^{*}$ cases shown in Fig. 3a (dots and diamonds) suggest that in the presence of significant remnant Coulomb repulsion ($U^{*}=U/3,U/2$ ) between paired electrons, large superconducting $T_{c}$ can be still maintained. In Fig. 3b and Fig. 3c, $T_{c}$ as a function of $U^{*}$ are shown at $n=0.6,0.7$ respectively (for fixed $U=8t$), where finite $T_{c}$ can be found even when $U^{*}\geq U/2$. Specifically, for the $n=0.6$ case, an enhanced $T_{c}\approx 0.05t$ is found at $U=8t,U^{*}=4t$. This means that in the large screening length limit $\lambda_{TF}\rightarrow\infty$ along $c$-axis, i.e., when Coulomb potential in the vertical direction becomes $U(r_{z})\propto\frac{e^{-r_{z}/\lambda_{TF}}}{r_{z}}\sim\frac{1}{r_{z}}$, (thus $U^{*}\sim U/2$ if layer distances $d_{12}=d_{23}$), the strong-coupling KL SC can still survive. This finding lie beyond the expectation of high-temperature DQMC results presented in Fig. 1d, where $V^{*}\sim-U/2$ for large $U$. In a purely static picture, $U^{*}\gtrsim U/2$ will lead to a net repulsion between top- and bottom-layer electrons [ i.e., $(|V^{*}|-U^{*})\lesssim 0$], thereby severely inhibiting pairing between them. We attribute the persistence of high $T_{c}$ under $U^{*}\gtrsim U/2$ condition to the retarded pairing or suppression of layer density fluctuations (see discussions in End Matter). A further remarkable point to emphasize is the sizeable $T_{c}$ up to a filling factor of $n=0.7$ (40% electron doping), as shown in Fig. 3c. This presents a clear difference with cases in Hubbard model for cuprates [15], where SC rapidly vanishes beyond $p\lesssim 30\%$ for hole-doping and $p\lesssim 20\%$ for electron-doping. This distinctness is rooted in the “pairing glue”: SC in cuprates relies on the collective spin fluctuations [12], which are suppressed at large dopings. In contrast, the strong-coupling KL SC identified here does not involve any intermediate collective modes, thus ensuring robustness across a broad range of $n$.
@@
The phase diagrams - Fig. 4a presents a superconducting phase diagram showing $T_{c}$ as a function of electron density $n$ for the spinless model on 2D triangular lattice from nearly empty ($n\sim 0.2$) to nearly full-filled ($n\sim 0.9$) at $U=6t,U^{*}=0$. One sees that $T_{c}$ is maximumized as $n$ being close to half-filling. Again, here we find SC instability in an extremely large range of $n$. Increasing repulsion $U$ can even further expand the SC regime. For instance, on the electron doped side, finite $T_{c}$ can be found for $n_{c}\gtrsim 0.95$ at $U=12t$ ( Fig. 4b).